# LITP-25-17

### v1

@@ 6.1 $\pi^{+}\pi^{0}$ Scattering
Eq. (6.5) is an all-order result in $\alpha_{\mathcal{E}}$. To unambiguously identify the relevant scales that enter the result, we can work in the perturbative regime $\alpha_{\mathcal{E}}\ll 1$ and include the full 1-loop $O(\alpha)$ effects. For illustrative reasons and simplicity, we will assume that loop effects induced by the Wilson coefficients – thoroughly studied in Refs. [Bellazzini:2020cot, Bellazzini:2021oaj, Beadle:2024hqg, Beadle:2025cdx, Chang:2025cxc, Peng:2025klv] – are small compared to $\alpha_{\mathcal{E}}$: $\alpha\ll g_{*}^{2}\ll\alpha_{\mathcal{E}}$, where $g_{*}^{2}\sim c_{i,j}M^{2i}$ denotes the generic EFT coupling strength.
@@
$c_{20}>0\,,\qquad(M^{2}-m_{\pm}^{2})\frac{c_{31}}{c_{20}}\geq-\frac{3}{2}\,.$ (6.9)
The smearing functions $\psi$ associated to these bounds satisfy simultaneously (5.13), as well as orthogonality to higher Wilson coefficients, via $\int_{\mathcal{E}}dq\psi(q)q^{2n}=0$ for $n\geq 2$, in order for the bound to be independent from them. For $n\to\infty$ such functionals tend to linear combinations of $\delta(q)$ and $\delta^{\prime}(q)$ as $\mathcal{E}\to 0$ [Beadle:2024hqg], and the problem becomes equivalent to matching $q$-derivatives in UV and IR.
@@
The constraints (6.11) are continuously connected to the constraints that hold with no electromagnetic interactions as long as $m^{2}_{\pm}$ remains finite in the limit $\alpha\to 0$. It is interesting to consider also the scenario where $m_{\pm}$ is not fixed, but $m_{\pm}^{2}\propto M^{2}\alpha/4\pi$, corresponding to a situation where electromagnetic interactions are the only source of explicit symmetry breaking. In this scenario, the bounds appear to change discretely. This is not very surprising since the IR mass singularities are correlated with the coupling constant [^20] A similar phenomenon indeed happens in the EFT of spin–3/2 particles of mass $m_{3/2}$, which is found inconsistent with positivity bounds whenever $M\gg m_{3/2}$, unless gravitons are also present and couple in a supersymmetric way [Bellazzini:2025shd] even in the decoupling limit $G\to 0$, since $m_{3/2}^{2}\propto G$. $1/m_{\pm}^{2}\sim 1/\alpha$; however, in this massless limit, collinear singularities – as manifest by the branch point of $\mathcal{W}^{+0}(t)$ at $4m_{\pm}^{2}\to 0$ – would become important and our calculation would have to be rethought. Nevertheless, as long as one restricts to a detector resolution $\mathcal{E}$ as small as $\sqrt{\alpha}M\mathrm{exp}(-\alpha_{\mathcal{E}}/\alpha)$ with $\alpha_{\mathcal{E}}\ll 1$, it’s still possible to perform a perturbative calculation within the framework developed here. It then follows again that $c_{31}M^{2}/c_{20}\geq-O(1)$.
@@
The previous analysis relied on a generic parametrization of the low-energy amplitude $\mathcal{M}_{+0-0}^{\mathrm{tree}}$, in which Wilson coefficients could take any value prior to the positivity bounds – this is how tree-level bounds are traditionally derived. This (unrealistic) setup becomes untreatable beyond tree-level, where loops of very irrelevant couplings contribute also to the lower arcs, and one has to assume they be smaller than the tree level coefficients $|c_{n,k}|\,M^{2n}/16\pi^{2}\lesssim 1$ for the ansatz to be consistently calculable [Bellazzini:2020cot]. In our specific case, the problem of hypotheses-less generic amplitudes is that they include also those in which the coefficients $c_{n,k}$ (controlled by UV physics at the scale $M$) completely undo the radiative IR effects of electromagnetism, associated with the scale $m_{\pm}$.
@@
If we neglect all terms larger than $N=4$, bounds of this type can be obtained semianalytically by directly studying the $q$-integrand of arcs as a function of $q$. The UV-representation of arcs implies that the inelastic $q\neq 0$ part is bounded by the elastic $q=0$ one, at the level of the $q$-integrand – this follows simply because Legendre polynomials are bounded by $|P_{\ell}|\leq 1$. In the IR part, ignoring EFT terms with $n\geq N=4$ we have only $c_{2,0}$ and $c_{3,1}$, and we obtain a simple inequality [Bellazzini:2023nqj, Bellazzini:2025shd],
@@ 6.2 $\pi^{+}\pi^{+}$ Scattering
We now consider the scattering process $\pi^{+}\pi^{+}\to\pi^{+}\pi^{+}$ within the same EFT. As in the $\pi^{+}\pi^{0}$ case, we work with stripped amplitudes in order to construct a well-defined analytic function free of IR divergences. There is, however, one important qualitative difference with respect to $\pi^{+}\pi^{0}$ scattering: the presence of a $t\to 0$ singularity, which would naively obstruct the derivation of positivity bounds in $D=4$ [Caron-Huot:2021rmr, Henriksson:2022oeu, Caron-Huot:2022ugt]. [^21] The $1/t$ tree level pole does not enter twice-subtracted dispersion relations. At loop level, however, box-type diagrams involving only electromagnetic interactions generate both a singular behavior in $q$ and non-analyticity in $s$ thereby propagating the singularity to all $n$-subtracted dispersion relations in the schematic form of $(\log q)/q^{2}$ terms. As explained in Section 5, working with IR–finite stripped amplitudes automatically resolves this issue: the $1/t$ singularity gives rise only to a finite, calculable amount of negativity, as we explicitly show here for $\pi^{+}\pi^{+}$-scattering.
@@
Positivity bounds in this context are conveniently derived at real fixed $q^{2}=tu/s$, via $t$-$u$ crossing symmetric dispersion relations [Bellazzini:2025shd], which follow the structure of Section 5 but are also explicitly expanded upon in Appendix A.1. Their appeal lies in the fact that, at tree-level, only finitely many Wilson coefficients enter into any IR arc. For instance,
@@
By instead choosing functionals that project out only the $\bar{c}_{4,2}$ term, we get bounds on $\bar{c}_{3,1}$ in terms of $\bar{c}_{2,0}$. In Fig. 4 we show in red the bounds obtained through the standard numerical algorithm [Caron-Huot:2021rmr]. In addition, we can obtain simple bounds with an analytic dependence on $\alpha_{\mathcal{E}}$ by optimizing on the $\bar{c}_{2,0}$ term while keeping fixed the $\bar{c}_{3,1}$ term and scanning different values of the normalization $\psi^{\prime}(0)$. For instance,
@@
Figure 4: Left panel: one-loop bounds on $M^{6}\bar{c}_{3,1}/\alpha_{\mathcal{E}}^{2}$ and $M^{4}\bar{c}_{2,0}/\alpha_{\mathcal{E}}^{2}$. In red the one–loop allowed region. The dashed lines are obtained by optimizing only on the $\bar{c}_{2,0}$ term and fixing $\psi^{\prime}(0)$ to 1828018280 (blue), $4\cdot 10^{4}$ (orange), while keeping fixed $\int_{\mathcal{E}}^{q_{max}}dq\,\psi(q)q^{2}=\pm 1$, and respectively correspond to the upper and lower bound in Eq. (6.21). Right panel: one-loop upper bound on $M^{2}\bar{c}_{3,1}/\bar{c}_{2,0}$ as a function of $\alpha_{\mathcal{E}}^{2}/(M^{4}\bar{c}_{2,0})$. The dashed blue line corresponds to the upper bound in Eq. (6.21), while the black dashed line is the tree-level bound.
@@ 7 Positivity Bounds with Gravity
$\mathcal{M}^{\mathrm{tree,GR}}=8\pi G\left(\frac{tu}{s}+\frac{us}{t}+\frac{st}{u}\right)+\hat{c}_{2,0}(s^{2}+t^{2}+u^{2})+\hat{c}_{3,1}stu+\hat{c}_{4,0}(s^{2}+t^{2}+u^{2})^{2}+\ldots$ (7.1)
@@
$\begin{split}\mathcal{A}_{0}^{(0)}[\psi,\mathcal{M}^{\textrm{tree}}]_{\mathrm{GR}}&=\lim_{\begin{subarray}{c}G\to 0\\ \mathcal{E}\to 0\\ GM^{2}\log M/\mathcal{E}\to G_{\mathcal{E}}\end{subarray}}\int_{\mathcal{E}}^{q_{max}}dq\,\psi(q)\,\left[\frac{8\pi G}{q^{2}}+2\hat{c}_{2,0}+q^{2}\hat{c}_{3,1}\right]\,.\end{split}$ (7.3)
For ordinary IR–divergent amplitudes, a similar expression is found in the literature; in that case, even at tree-level, in $D=4$ dimensions, functionals that are positive in the UV representation are not integrable against the pole in $q$ as the IR–cutoff is removed [Caron-Huot:2021rmr]. Instead, for IR–finite stripped amplitudes, the amount of negativity in the integral in (7.3) is both finite and of order $O(GM^{2},\mathcal{E}^{2}/M^{2})$, see Section 5, and can therefore be discarded in the scaling limit (1.1).
At loop level, the $G_{\mathcal{E}}^{2}$ term contributing to the arcs is obtained from the first line of the full $G^{2}$ amplitude in Eq. (C.3). Following [Beadle:2025cdx, Chang:2025cxc], the arc with $n=0$ evaluates to
$\begin{split}\delta\mathcal{A}_{0}^{(0)}[\psi,\mathcal{M}_{\mathcal{E}}^{\mathrm{1-loop\,GR}}]&=\lim_{\begin{subarray}{c}G\to 0\\ \mathcal{E}\to 0\\ GM^{2}\log M/\mathcal{E}\to G_{\mathcal{E}}\end{subarray}}\int_{\mathcal{E}}^{q_{max}}dq\,\psi(q)\,\frac{16G^{2}M^{2}}{q^{2}}\log(\frac{\mathcal{E}^{2}}{M^{2}})\,.\end{split}$ (7.4)
@@
while releasing this constraint we obtain two–sided bounds on $\hat{c}_{3,1}q^{2}$, following the standard procedure [Caron-Huot:2021rmr, Caron-Huot:2022ugt]. Analogously to the $\pi_{+}\pi_{+}$ case, we can keep the $G_{\mathcal{E}}$ dependence analytic and derive constraints of the form
@@
respectively obtained with $\psi^{\prime}(0)=350$ and $\psi^{\prime}(0)=630$. The bounds are shown in Fig. 5. As in the $\pi_{+}\pi_{+}$ case, we did not focus on optimizing the constraints. Importantly, as $G_{\mathcal{E}}$ goes to zero we recover the tree-level bounds in absence of gravity $\hat{c}_{2,0}>0$, $\hat{c}_{3,1}M^{2}<\frac{3}{2}\hat{c}_{2,0}$, as highlighted in Fig. 6. Also the lower bound on $\hat{c}_{3,1}$ is consistent with the analytic tree–level bound [Caron-Huot:2021rmr].
@@ 8 Summary and Conclusions
To overcome this obstruction, we introduced stripped amplitudes $\mathcal{M}_{\mathcal{E}}$, a one–parameter family of functions obtained by factoring out from standard amplitudes the analytic continuation of the Weinberg soft exponential. For any $\mathcal{E}>0$, the resulting $\mathcal{M}_{\mathcal{E}}$ are IR finite, analytic, crossing symmetric, and Lorentz invariant. For $\mathcal{E}$ exponentially small relative to any other hard scale $M$, this parameter acquires a clear physical meaning: it coincides with the energy resolution $\Delta E=\mathcal{E}$ of a macroscopic device capable of detecting soft photons or gravitons (2.6). Different values of $\mathcal{E}$ correspond to different experimental apparatuses.
@@
Relation to previous literature. Our construction differs qualitatively from earlier attempts to incorporate long–range forces into positivity bounds. Previous analyses established nontrivial gravitational bounds only for $D>4$ [Caron-Huot:2021rmr] or in AdS spacetime [Caron-Huot:2021enk], with the bounds disappearing as $D\to 4$ or the curvature is removed. In flat spacetime, working with IR–divergent amplitudes at fixed order in $G$ ($\alpha$) obstructs the existence of positive functionals: they must be integrable against the $1/t$ singularity, which fails in $D=4$. Introducing hard IR cutoffs by hand [Caron-Huot:2021rmr, Henriksson:2022oeu, Caron-Huot:2022ugt] does not resolve the issue: either the IR part of the dispersion relation diverges or the UV part becomes unconstrained. Moreover, the unitarity and analytic properties of the regulated amplitudes remain unclear in this case, so as their meaning. Furthermore, as the IR cutoff is removed, the IR logarithms become large and require resummation of all orders in $G$ ($\alpha$), as emphasized in [Chang:2025cxc], giving rise to apparently paradoxical non–decoupling effects, as we discuss below. Other approaches focusing solely on the graviton pole [Hamada:2018dde, Bellazzini:2019xts, Tokuda:2020mlf, Herrero-Valea:2020wxz, Noumi:2021uuv, Noumi:2022wwf] address only one aspect of the IR problem.
@@
In contrast, our framework removes the universal IR divergences from the outset by working with the IR–finite $\mathcal{M}_{\mathcal{E}}$, where the IR regulator $\epsilon$ has already been taken to zero. The remaining scale, when exponentially smaller than all other hard scales, acquires a physical meaning in terms of detector resolution, see (2.6), and should not be taken to zero. In the detector scaling limit, potentially negative contributions remain finite and we are able to suppress them, thanks to the small parameter $G/G_{\mathcal{E}}=1/\log(M/\mathcal{E})\ll 1$. This logic is fully analogous to the $1/N^{2}$ parametric suppression familiar from large–$N$ positivity bounds [Albert:2022oes, Fernandez:2022kzi, Albert:2023jtd, Ma:2023vgc, Dong:2024omo], in which loops, specific uu–channel contributions, and arcs at infinity in once–subtracted dispersions are systematically neglected. In our setting this is achieved by changing an experimental parameter $\mathcal{E}$ rather than $N$ and the theory itself. Moreover, while $G_{\mathcal{E}}/G\gg 1$, we can chose $\mathcal{E}$ such that $G_{\mathcal{E}}$ is perturbative, avoiding the need to resum large logarithms which would effectively reintroduce IR divergences.
@@
Bounds derived in AdS space [Caron-Huot:2021rmr] are conceptually closest to ours, in that the AdS curvature radius may be tentatively interpreted as an effective resolution scale $1/\mathcal{E}$, rather than being taken to zero, although a precise recasting of AdS–bounds at finite radius in terms of flat space quantities has not been derived in the literature. By contrast, extrapolation of $D>4$ bounds to $D\to 4$ by holding $GM^{2}/(D-4)$ fixed, would provide technically finite results, but lack a physically meaningful and quantitative interpretation.
@@
We also comment on the very interesting analysis of [Chang:2025cxc], which highlighted the necessity of resumming radiative effects in the vicinity of the $t=0$ region for the full amplitude. Indeed, fixed–order expansions in $GM^{2}/\epsilon$ are not meaningful once $\epsilon\to 0$, since the effective expansion parameter becomes large independently of how small $GM^{2}$ is. Ref. [Chang:2025cxc] further showed that, when the limits are taken in the correct order, the $1/t$ pole is not by itself the fundamental obstruction to positivity bounds in $D=4$, as one can construct positive functionals that integrate the purely gravitational contribution to the dispersion relations to finite values. For a similar perspective see also [FuentesZamoro:2025exp].
@@
At the same time, the analysis of [Chang:2025cxc] led to apparently GG–independent bounds—for example, $g_{2}M^{4}+259\geq 0$ for amplitudes of the form $\mathcal{M}=g_{2}(s^{2}+t^{2}+u^{2})+\cdots$. This behavior suggested a puzzling lack of decoupling in the limit $G\to 0$, since the unperturbed theory instead satisfies $g_{2}\geq 0$. We resolve this puzzle by observing that IR divergences suppress the full amplitude at all kinematic scales, not only in the neighborhood of $t=0$ where the $1/t$ contribution from graviton exchange arises. As a consequence, bounds derived from IR–divergent amplitudes can formally be written, but effectively reduce to trivial statements of the type $g_{2}M^{4}\times 0+259\geq 0$. This should be contrasted with our use of IR–finite stripped amplitudes, which avoid this pathology altogether and yield finite bounds with a smooth $G\to 0$ limit, see Section 5 and explicit examples in Section 7. Once IR divergences have been removed, and only then, there is no need to work with the “optimal resolution” detector to efficiently constrain irrelevant Wilson coefficients.

### v2

@@ 6.1 $\pi^{+}\pi^{0}$ Scattering
Eq. (6.5) is an all-order result in $\alpha_{\mathcal{E}}$. To unambiguously identify the relevant scales that enter the result, we can work in the perturbative regime $\alpha_{\mathcal{E}}\ll 1$ and include the full 1-loop $O(\alpha)$ effects. For illustrative reasons and simplicity, we will assume that loop effects induced by the Wilson coefficients – thoroughly studied in Refs. [19, 26, 16, 17, 53, 125] – are small compared to $\alpha_{\mathcal{E}}$: $\alpha\ll g_{*}^{2}\ll\alpha_{\mathcal{E}}$, where $g_{*}^{2}\sim c_{i,j}M^{2i}$ denotes the generic EFT coupling strength.
@@
$c_{2,0}>0\,,\qquad(M^{2}-m_{\pm}^{2})\frac{c_{3,1}}{c_{2,0}}\geq-\frac{3}{2}\,.$ (6.9)
The smearing functions $\psi$ associated to these bounds satisfy simultaneously (5.13), as well as orthogonality to higher Wilson coefficients, via $\int_{\mathcal{E}}dq\psi(q)q^{2n}=0$ for $n\geq 2$, in order for the bound to be independent from them. For $n\to\infty$ such functionals tend to linear combinations of $\delta(q)$ and $\delta^{\prime}(q)$ as $\mathcal{E}\to 0$ [16], and the problem becomes equivalent to matching $q$-derivatives in UV and IR.
@@
The constraints (6.11) are continuously connected to the constraints that hold with no electromagnetic interactions as long as $m^{2}_{\pm}$ remains finite in the limit $\alpha\to 0$. It is interesting to consider also the scenario where $m_{\pm}$ is not fixed, but $m_{\pm}^{2}\propto M^{2}\alpha/4\pi$, corresponding to a situation where electromagnetic interactions are the only source of explicit symmetry breaking. In this scenario, the bounds appear to change discretely. This is not very surprising since the IR mass singularities are correlated with the coupling constant [^22] A similar phenomenon indeed happens in the EFT of spin–3/2 particles of mass $m_{3/2}$, which is found inconsistent with positivity bounds whenever $M\gg m_{3/2}$, unless gravitons are also present and couple in a supersymmetric way [25] even in the decoupling limit $G\to 0$, since $m_{3/2}^{2}\propto G$. $1/m_{\pm}^{2}\sim 1/\alpha$; however, in this massless limit, collinear singularities – as manifest by the branch point of $\mathcal{W}^{+0}(t)$ at $4m_{\pm}^{2}\to 0$ – would become important and our calculation would have to be rethought. Nevertheless, as long as one restricts to a detector resolution $\mathcal{E}$ as small as $\sqrt{\alpha}M\mathrm{exp}(-\alpha_{\mathcal{E}}/\alpha)$ with $\alpha_{\mathcal{E}}\ll 1$, it’s still possible to perform a perturbative calculation within the framework developed here. It then follows again that $c_{3,1}M^{2}/c_{2,0}\geq-O(1)$.
@@
The previous analysis relied on a generic parametrization of the low-energy amplitude $\mathcal{M}_{+0-0}^{\mathrm{tree}}$, in which Wilson coefficients could take any value prior to the positivity bounds – this is how tree-level bounds are traditionally derived. This (unrealistic) setup becomes untreatable beyond tree-level, where loops of very irrelevant couplings contribute also to the lower arcs, and one has to assume they be smaller than the tree level coefficients $|c_{n,k}|\,M^{2n}/16\pi^{2}\lesssim 1$ for the ansatz to be consistently calculable [19]. In our specific case, the problem of hypotheses-less generic amplitudes is that they include also those in which the coefficients $c_{n,k}$ (controlled by UV physics at the scale $M$) completely undo the radiative IR effects of electromagnetism, associated with the scale $m_{\pm}$.
@@
If we neglect all terms larger than $N=4$, bounds of this type can be obtained semianalytically by directly studying the $q$-integrand of arcs as a function of $q$. The UV-representation of arcs implies that the inelastic $q\neq 0$ part is bounded by the elastic $q=0$ one, at the level of the $q$-integrand – this follows simply because Legendre polynomials are bounded by $|P_{\ell}|\leq 1$. In the IR part, ignoring EFT terms with $n\geq N=4$ we have only $c_{2,0}$ and $c_{3,1}$, and we obtain a simple inequality [21, 25],
@@ 6.2 $\pi^{+}\pi^{+}$ Scattering
We now consider the scattering process $\pi^{+}\pi^{+}\to\pi^{+}\pi^{+}$ within the same EFT. As in the $\pi^{+}\pi^{0}$ case, we work with stripped amplitudes in order to construct a well-defined analytic function free of IR divergences. There is, however, one important qualitative difference with respect to $\pi^{+}\pi^{0}$ scattering: the presence of a $t\to 0$ singularity, which would naively obstruct the derivation of positivity bounds in $D=4$ [49, 98, 46]. [^23] The $1/t$ tree level pole does not enter twice-subtracted dispersion relations. At loop level, however, box-type diagrams involving only electromagnetic interactions generate both a singular behavior in $q$ and non-analyticity in $s$ thereby propagating the singularity to all $n$-subtracted dispersion relations in the schematic form of $(\log q)/q^{2}$ terms. As explained in Section 5, working with IR–finite stripped amplitudes automatically resolves this issue: the $1/t$ singularity gives rise only to a finite, calculable amount of negativity, as we explicitly show here for $\pi^{+}\pi^{+}$-scattering.
@@
Positivity bounds in this context are conveniently derived at real fixed $q^{2}=tu/s$, via $t$-$u$ crossing symmetric dispersion relations [25], which follow the structure of Section 5 but are also explicitly expanded upon in Appendix A.1. Their appeal lies in the fact that, at tree-level, only finitely many Wilson coefficients enter into any IR arc. For instance,
@@
By instead choosing functionals that project out only the $\bar{c}_{4,2}$ term, we get bounds on $\bar{c}_{3,1}$ in terms of $\bar{c}_{2,0}$. In Fig. 4 we show in red the bounds obtained through the standard numerical algorithm [49]. In addition, we can obtain simple bounds with an analytic dependence on $\alpha_{\mathcal{E}}$ by optimizing on the $\bar{c}_{2,0}$ term while keeping fixed the $\bar{c}_{3,1}$ term and scanning different values of the normalization $\psi^{\prime}(0)$. For instance,
@@
Figure 4: Left panel: one-loop bounds on $M^{6}\bar{c}_{3,1}/\alpha_{\mathcal{E}}^{2}$ and $M^{4}\bar{c}_{2,0}/\alpha_{\mathcal{E}}^{2}$. In red the one–loop allowed region. The dashed lines are obtained by optimizing only on the $\bar{c}_{2,0}$ term and fixing $\psi^{\prime}(0)$ to $18280$ (blue), $4\cdot 10^{4}$ (orange), while keeping fixed $\int_{\mathcal{E}}^{q_{max}}dq\,\psi(q)q^{2}=\pm 1$, and respectively correspond to the upper and lower bound in Eq. (6.21). Right panel: one-loop upper bound on $M^{2}\bar{c}_{3,1}/\bar{c}_{2,0}$ as a function of $\alpha_{\mathcal{E}}^{2}/(M^{4}\bar{c}_{2,0})$. The dashed blue line corresponds to the upper bound in Eq. (6.21), while the black dashed line is the tree-level bound.
@@ 7 Positivity Bounds with Gravity
$\mathcal{M}^{\mathrm{tree,GR}}=8\pi G\left(\frac{tu}{s}+\frac{us}{t}+\frac{st}{u}\right)+\hat{c}_{2,0}\left(\!\frac{s^{2}+t^{2}+u^{2}}{2}\!\right)+\hat{c}_{3,1}stu+\hat{c}_{4,0}\left(\!\frac{s^{2}+t^{2}+u^{2}}{2}\!\right)^{2}\!+\ldots$ (7.1)
@@
$\begin{split}\mathcal{A}_{0}^{(0)}[\psi,\mathcal{M}^{\textrm{tree,\,GR}}]&=\lim_{\begin{subarray}{c}G\to 0\\ \mathcal{E}\to 0\\ GM^{2}\log M/\mathcal{E}\to G_{\mathcal{E}}\end{subarray}}\int_{\mathcal{E}}^{q_{max}}dq\,\psi(q)\,\left[\frac{8\pi G}{q^{2}}+\hat{c}_{2,0}+q^{2}\hat{c}_{3,1}\right]\,.\end{split}$ (7.3)
For ordinary IR–divergent amplitudes, a similar expression is found in the literature; in that case, even at tree-level, in $D=4$ dimensions, functionals that are positive in the UV representation are not integrable against the pole in $q$ as the IR–cutoff is removed [49]. Instead, for IR–finite stripped amplitudes, the amount of negativity in the integral in (7.3) is both finite and of order $O(GM^{2},\mathcal{E}^{2}/M^{2})$, see Section 5, and can therefore be discarded in the scaling limit (1.1).
At loop level, the $G_{\mathcal{E}}^{2}$ term contributing to the arcs is obtained from the first line of the full $G^{2}$ amplitude in Eq. (C.3). Following [17, 53], the arc with $n=0$ evaluates to
$\begin{split}\delta\mathcal{A}_{0}^{(0)}[\psi,\mathcal{M}_{\mathcal{E}}^{\mathrm{1-loop,\,GR}}]&=\lim_{\begin{subarray}{c}G\to 0\\ \mathcal{E}\to 0\\ GM^{2}\log M/\mathcal{E}\to G_{\mathcal{E}}\end{subarray}}\int_{\mathcal{E}}^{q_{max}}dq\,\psi(q)\,\frac{16G^{2}M^{2}}{q^{2}}\log(\frac{\mathcal{E}^{2}}{M^{2}})\,.\end{split}$ (7.4)
@@
while releasing this constraint we obtain two–sided bounds on $\hat{c}_{3,1}q^{2}$, following the standard procedure [49, 46]. Analogously to the $\pi_{+}\pi_{+}$ case, we can keep the $G_{\mathcal{E}}$ dependence analytic and derive constraints of the form
@@
respectively obtained with $\psi^{\prime}(0)=350$ and $\psi^{\prime}(0)=630$. The bounds are shown in Fig. 5. As in the $\pi_{+}\pi_{+}$ case, we did not focus on optimizing the constraints. Importantly, as $G_{\mathcal{E}}$ goes to zero we recover the tree-level bounds in absence of gravity $\hat{c}_{2,0}>0$, $\hat{c}_{3,1}M^{2}<\frac{3}{2}\hat{c}_{2,0}$, as highlighted in Fig. 6. Also the lower bound on $\hat{c}_{3,1}$ is consistent with the analytic tree–level bound [49].
@@ 8 Summary and Conclusions
To overcome this obstruction, we introduced stripped amplitudes $\mathcal{M}_{\mathcal{E}}$, a one–parameter family of functions obtained by factoring out from standard amplitudes the analytic continuation of the Weinberg soft exponential. For any $\mathcal{E}>0$, the resulting $\mathcal{M}_{\mathcal{E}}$ are (exactly) IR finite, analytic, crossing symmetric, and Lorentz invariant. For $\mathcal{E}$ exponentially small relative to any other hard scale $M$, this parameter acquires a clear physical meaning: it coincides with the energy resolution $\Delta E=\mathcal{E}$ of a macroscopic device capable of detecting soft photons or gravitons (2.6). Different values of $\mathcal{E}$ correspond to different experimental apparatuses.
@@
Remarkably, subleading tree-level $O(\alpha,G)$ contributions can also be controlled quantitatively, provided one assumes more about the UV completion. If the UV remains in a tree-level regime—namely, the high-energy amplitude is well approximated by a meromorphic function, as in large-$N$ theories [6, 79, 34] or weakly-coupled stringy gravitational completions [44]—then the residual error on the bounds is reduced to genuine loop effects, schematically $O(\alpha^{2},G^{2},\alpha g_{*}^{2},Gg_{*}^{2},\alpha\alpha_{\mathcal{E}},\ldots)$, while the tree-level $O(\alpha,G)$ terms remain under control. This is stronger than simply assuming a weakly-coupled UV completion, which may still contribute through loops. It is really an assumption about the leading analytic structure of the amplitudes. Moreover, such a statement only makes sense for IR-safe observables, such as the stripped amplitude $\mathcal{M}_{\mathcal{E}}$, since otherwise loop corrections would invalidate the tree-level approximation for the full amplitude $\mathcal{M}$ irrespective of the smallness of the couplings.
@@
Relation to previous literature. Our construction differs qualitatively from earlier attempts to incorporate long–range forces into positivity bounds. Previous analyses established nontrivial gravitational bounds only for $D>4$ [49] or in AdS spacetime [48], with the bounds disappearing as $D\to 4$ or the curvature is removed. In flat spacetime, working with IR–divergent amplitudes at fixed order in $G$ ($\alpha$) obstructs the existence of positive functionals: they must be integrable against the $1/t$ singularity, which fails in $D=4$. Introducing hard IR cutoffs by hand [49, 98, 46] does not resolve the issue: either the IR part of the dispersion relation diverges or the UV part becomes unconstrained. Moreover, the unitarity and analytic properties of the regulated amplitudes remain unclear in this case, so as their meaning. Furthermore, as the IR cutoff is removed, the IR logarithms become large and require resummation of all orders in $G$ ($\alpha$), as emphasized in [53], giving rise to apparently paradoxical non–decoupling effects, as we discuss below. Other approaches focusing solely on the graviton pole [90, 23, 136, 99, 119, 120] address only one aspect of the IR problem.
@@
In contrast, our framework removes the universal IR divergences from the outset by working with the IR–finite $\mathcal{M}_{\mathcal{E}}$, where the IR regulator $\epsilon$ has already been taken to zero. The remaining scale, when exponentially smaller than all other hard scales, acquires a physical meaning in terms of detector resolution, see (2.6), and should not be taken to zero. In the detector scaling limit, potentially negative contributions remain finite and we are able to suppress them, thanks to the small parameter $G/G_{\mathcal{E}}=1/\log(M/\mathcal{E})\ll 1$. This logic is fully analogous to the $1/N^{2}$ parametric suppression familiar from large–$N$ positivity bounds [6, 79, 7, 113, 70], in which loops, specific uu–channel contributions, and arcs at infinity in once–subtracted dispersions are systematically neglected. In our setting this is achieved by changing an experimental parameter $\mathcal{E}$ rather than $N$ and the theory itself. Moreover, while $G_{\mathcal{E}}/G\gg 1$, we can chose $\mathcal{E}$ such that $G_{\mathcal{E}}$ is perturbative, avoiding the need to resum large logarithms which would effectively reintroduce IR divergences.
@@
Bounds derived in AdS space [49] are conceptually closest to ours, in that the AdS curvature radius may be tentatively interpreted as an effective resolution scale $1/\mathcal{E}$, rather than being taken to zero, although a precise recasting of AdS–bounds at finite radius in terms of flat space quantities has not been derived in the literature. By contrast, extrapolation of $D>4$ bounds to $D\to 4$ by holding $GM^{2}/(D-4)$ fixed, would provide technically finite results, but lack a physically meaningful and quantitative interpretation.
@@
We also comment on the very interesting analysis of [53], which highlighted the necessity of resumming radiative effects in the vicinity of the $t=0$ region for the full amplitude. Indeed, fixed–order expansions in $GM^{2}/\epsilon$ are not meaningful once $\epsilon\to 0$, since the effective expansion parameter becomes large independently of how small $GM^{2}$ is. Ref. [53] further showed that, when the limits are taken in the correct order, the $1/t$ pole is not by itself the fundamental obstruction to positivity bounds in $D=4$, as one can construct positive functionals that integrate the purely gravitational contribution to the dispersion relations to finite values. For a similar perspective see also [40, 82].
@@
At the same time, the analysis of [53] led to apparently GG–independent bounds—for example, $g_{2}M^{4}+259\geq 0$ for amplitudes of the form $\mathcal{M}=g_{2}(s^{2}+t^{2}+u^{2})+\cdots$. This behavior suggested a puzzling lack of decoupling in the limit $G\to 0$, since the unperturbed theory instead satisfies $g_{2}\geq 0$. We resolve this puzzle by observing that IR divergences suppress the full amplitude at all kinematic scales, not only in the neighborhood of $t=0$ where the $1/t$ contribution from graviton exchange arises. As a consequence, bounds derived from IR–divergent amplitudes can formally be written, but effectively reduce to trivial statements of the type $g_{2}M^{4}\times 0+259\geq 0$. This should be contrasted with our use of IR–finite stripped amplitudes, which avoid this pathology altogether and yield finite bounds with a smooth $G\to 0$ limit, see Section 5 and explicit examples in Section 7. Once IR divergences have been removed, and only then, there is no need to work with the “optimal resolution” detector to efficiently constrain irrelevant Wilson coefficients.